PUMA: An Introduction

• Parallel Unstructured Maritime Aerodynamics.

• Computer program for analysis of internal and external nonreacting compressible flows over arbitrarily complex 3D geometries.

• Written entirely in ANSI C using MPI library for message passing and hence highly portable giving good performance.

• Based on Finite Volume method and supports mixed topology unstructured grids composed of tetrahedra, wedges, pyramids and hexahedra (bricks).

• May be run so as to preserve time accuracy or as a pseudounsteady formulation to enhance convergence to steady-state.

• Uses dynamic memory allocation, thus problem size is limited only by the amount of memory available on the machine.

• Needs 582 bytes/cell and 634 bytes/face using double precision variables (not including message passing overhead).

• The development of PUMA was highly motivated by the lack of work that compares the performance of several different algorithms on the same problem. Hence PUMA implements a range of time-integration schemes like Runge-Kutta, Jacobi and various Successive Over-relaxation Schemes (SOR), as well as both Roe and Van Leer numerical flux schemes. It also implements various monotone limiters used in secondorder computations (Venkatakrishnan, Barth, Van Albada, Superbee).